OpenStudy (anonymous):

if y varies inversely with x, and y=-16 when x=-64, what is the constant of variation?

OpenStudy (anonymous):

y = kx is the formula

OpenStudy (anonymous):

y = kx -16 = -64x k = 0.25

OpenStudy (anonymous):

* -16 = -64k

OpenStudy (anonymous):

So the answer is 64? @greenlegodude57

OpenStudy (anonymous):

No, it's 0.25

OpenStudy (anonymous):

Ugh I hate math @greenlegodude57 thank you

OpenStudy (anonymous):

xD

OpenStudy (johnweldon1993):

Not quite... Inversely means the formula would be \[\large y = \frac{k}{x}\] \[\large -16 = \frac{k}{-64}\] \[\large k = -16 \times -64 = 1024\]

OpenStudy (johnweldon1993):

Direct Variation \(\large y = kx\) Varying Inversely \(\large y = \frac{k}{x}\)

OpenStudy (anonymous):

Oh..

OpenStudy (anonymous):

My Bad.